OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (2, -9) and (4, 1). Please show all of your work.

OpenStudy (bahrom7893):

kiki it's time u learned this

OpenStudy (bahrom7893):

i solved like 5 of these for u.

OpenStudy (bahrom7893):

with solutions and steps.

OpenStudy (bahrom7893):

M = (y2-y1)/(x2-x1)

OpenStudy (bahrom7893):

u are given two points: (x1;y1) and (x2;y2). Plug them in. Easy.

OpenStudy (katrinakaif):

Find the slope of the two given points using : y2-y1 ----- x2-x1 Your points: (2,-9) is equated with (x1,y1) and your points (4,1) are in subtituted with (x2,y2) Solve this, and then plug it in as 'm' in the equation: (y-y1) = m(x-x1) Choose any given coordinate points to substitute for (x1,y1) and Evaluate!